OpenStudy (anonymous):

What is the solution set of |x – 6| = 10?

OpenStudy (anonymous):

(x-6)=10 or -(x-6)=10

OpenStudy (anonymous):

this -(x-6)=10 turns to (x-6)=-10

OpenStudy (anonymous):

|x-6|=10 Remove the absolute value term. This creates a \ on the right-hand side of the equation because |x|=\x. x-6=\(10) Set up the portion of the \ solution. x-6=10 Move all terms not containing x to the right-hand side of the equation. x=16 Set up the - portion of the \ solution. x-6=-(10) Multiply -1 by the 10 inside the parentheses. x-6=-10 Since -6 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 6 to both sides. x=6-10 Subtract 10 from 6 to get -4. x=-4 The solution to the equation includes both the positive and negative portions of the solution. x=16,-4

OpenStudy (anonymous):

\[|x-6| = 10 \implies x - 6 = 10 \ \ \text{ or }\ \ x - 6 = -10\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

ok any more questions ?